OpenStudy (anonymous):

Variables in Combinations? How would I turn C(x, 10) into an algebraic expression?

OpenStudy (anonymous):

\[_nC_k=\frac{n!}{k!(n-k)!}\] so \[_xC_{10}=\frac{x!}{10!(x-10)!}\]

OpenStudy (anonymous):

^^ can also be further reduced. \[x!=x(x-1)(x-2)\cdots(x-10)!\] so \[{}_xC_{10}=\frac{x!}{10!(x-10)!}=\frac{x(x-1)(x-2)\cdots(x-10)!}{10!(x-10)!}=\frac{x(x-1)(x-2)\cdots(x-9)}{10!}\]

OpenStudy (anonymous):

It can actually be found a different way.... I figured it out